OpenStudy (anonymous):

Consider the function f(x) = x2 + 3x − 7. Find the value of f(x) when x = 3 and select the correct answer below. 5 8 11 18

OpenStudy (anonymous):

You are given the function: \[f(x) = x^2 + 3x - 7\] Input \(f(3) \Rightarrow\) into your given function. \[f(3) = 3^2 + 3(3) - 7\]

OpenStudy (anonymous):

Do you know where to go from there?

OpenStudy (anonymous):

would it be f(3)=9+9-7 ?

OpenStudy (anonymous):

then it would equal 11 right ?

OpenStudy (anonymous):

Correct. Good job.

OpenStudy (anonymous):

thank you and could u help with one more

OpenStudy (anonymous):

Which of the following expressions is the inverse of the function y equals quantity x minus 2 divided by 3? y = 3x + 2 y = 2x + 3 y equals quantity x minus 3 divided by 2 y equals quantity x plus 2 divided by 3

OpenStudy (anonymous):

Sorry, I do not understand that question. I would help if I did.